OpenStudy (anonymous):

Which function has a removable discontinuity? A. p(x)=x/x^2-x-2 B. g(x)=2x-1/x C. f(x)=5x/1-x^2 D. h(x)=x^2-x-2/x+1

OpenStudy (shamil98):

Removable discontinuity means that the top and bottom have a value of that sets the fraction to 0/0.

OpenStudy (anonymous):

so how can i figure out which one has a removable discontinuity?

OpenStudy (shamil98):

Find the domain of the function. and Plug in that domain for x.

OpenStudy (shamil98):

the domain of a rational function is found by setting the denominator to equal 0

OpenStudy (shamil98):

h(x)=x^2-x-2/x+1 x + 1 = 0 x= -1 =(-1)^2 + 1 - 2/ -1 + 1 =0/0

OpenStudy (anonymous):

so does that mean that D would be the one with a removable discontinuity?

OpenStudy (shamil98):

yeah

OpenStudy (shamil98):

\[[x \in \mathbb{R}: x \ne - 1]\] is your domain btw.

OpenStudy (anonymous):

I see, so for this problem would i do the same thing and set everything to 0?

OpenStudy (anonymous):

why is that?

OpenStudy (shamil98):

well actually im not sure.. x-3^x / x^2 or it might be 2/x ..

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

is there a way to figure out which one it is for sure?

OpenStudy (shamil98):

Probably, bear with me i'm on three hours of sleep.. one sec..

OpenStudy (anonymous):

hahah of course, thank you :)

OpenStudy (shamil98):

Yeah, it's 2/x

OpenStudy (anonymous):

thank you soo much :)